OpenStudy (anonymous):

for the function f (x)=2x^2-10x+5 find f(0.5)

OpenStudy (anonymous):

f(0.5)=2(0.5)^2-10(0.5)+5 Then solve..

OpenStudy (anonymous):

is the answer 5?

OpenStudy (anonymous):

Did you use the order of operations?

OpenStudy (anonymous):

Yes, I just want to make sure I am right.

OpenStudy (anonymous):

Try again..

OpenStudy (anonymous):

f(0.5)=2(0.5)^2-10(0.5)+5 0.5^2=0.25 x 2=0.5 0.5-(10 x 0.5)=-4.5 -4.5+5= 0.5

OpenStudy (anonymous):

thank you so much.

OpenStudy (anonymous):

you're welcome!